The Haber process is used to synthesize ammonia (NH3) on a large scale by the reaction of nitrogen gas with hydrogen gas as follows:

N2 (g) + 3 H2 (g) → 2 NH3 (g)



Consider a reaction between a 7.89 L flask containing nitrogen gas and a 16.4 L flask of hydrogen gas. Both gases have a temperature of 295 K and the pressures inside both flasks is 0.86 bar.

What mass of ammonia (in g) would you expect to be produced at temperature 295 K and a pressure of 0.86 bar?
1
Expert's answer
2020-09-30T04:22:33-0400

Avogadro's law states: equal volumes of all gases, at the same temperature and pressure, have the same number of molecules.


It means that at the same temperature and pressure 7.89 L of nitrogen needs 7.89*3 = 23.67 L of hydrogen according to the equation. We have less hydrogen, so it is a limiting reagent. Find the volume of ammonia we can obtain from 16.4 L of hydrogen:

"V(NH_3) = 16.4*2\/3 = 10.9 L"


The combined gas law states:

"PV =mRT\/M", where R = gas constant (8.314*10-2 L*bar*K-1*mol-1), M -- molar mass (17 g/mol)

Find m:

"m = MPV\/RT = 17 * 0.86 * 10.9*100\/(8.314*295) = 6.5 (g)"

Answer: 6.5 g of ammonia would be expected to be produced at temperature 295 K and a pressure of 0.86 bar?